OpenStudy (anonymous):

I will award a medal!!!!!!!!!!!! help me please Find the circumference of a circle with radius of 18 in. Leave your answer in terms of Pi. 18 Pi. in. 324 Pi in. 54 Pi in. 36 Pi in. *Pi. stands for 3.14:)

OpenStudy (anonymous):

circumference = 2 x 3.14 x radius

OpenStudy (zzr0ck3r):

circumference = 2*pi*radius radius = 18 2*pi*18=2*18*pi=?

OpenStudy (anonymous):

Circumference = 2 • π • radius = π • diameter

OpenStudy (anonymous):

Circumference = 2 • π • radius Circumference = 2 • π • 18 Circumference = 36 • π